OpenStudy (anonymous):

2∫(0.01sen3πx)×(senπx)dx interval: 0 - 1

OpenStudy (anonymous):

the answer may not result in zero

OpenStudy (anonymous):

$$\cos(2\pi x)=\cos(3\pi x-\pi x)=\cos(3\pi x)\cos(\pi x)+\sin(3\pi x)\sin(\pi x)\\\cos(4\pi x)=\cos(3\pi x+\pi x)=\cos(3\pi x)\cos(\pi x)-\sin(3\pi x)\sin(\pi x)$$

OpenStudy (anonymous):

therefore \(\cos(2\pi x)-\cos(4\pi x)=2\sin(3\pi x)\sin(\pi x)\) allowing us to rewrite our integral:$$\begin{align*}0.01\int_0^1(\cos(2\pi x)-\cos(4\pi x))\ dx&=0.01\left(\frac1{2\pi}\sin(2\pi x)-\frac1{4\pi}\sin(4\pi x)\right)_0^1\\&=0.01\left(-\left(\frac1{2\pi}-\frac1{4\pi}\right)\right)\\&=-\frac1{400\pi}\end{align*}$$

OpenStudy (anonymous):

oops actually I forgot \(\sin0=0\ne 1\) so actually the second line should read \(0.01\cdot0\) and our integral result is just \(0\) -- cool!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we can use trigonometric identity

OpenStudy (anonymous):

but my result is zero

OpenStudy (anonymous):

this is because \(\sin(nx),\sin(mx)\) are orthogonal on \([0,\pi]\) for \(n,m\) both either odd or even

OpenStudy (anonymous):

and yes the result is \(0\) I wrote the wrong answer down above the integral evaluates to \(0\)

OpenStudy (anonymous):

but the result can not be zero

OpenStudy (anonymous):

but it is a wave equation, there must be an outcome

OpenStudy (anonymous):

2∫(0.01sen3πx)×(senπnx)dx interval: 0 - 1

OpenStudy (anonymous):

have the letter N

OpenStudy (anonymous):

okay if it has the letter \(n\) then the answer is \(0\) for all odd \(n=2k+1\) but the result is different for even \(n=2k\)

OpenStudy (anonymous):

actually, it's *always* \(0\) for integers \(n,m\)! http://www.wolframalpha.com/input/?i=int_0%5Epi+sin%28m+x%29+sin%28n+x%29+dx